### v2

# Unveiling Typographic Deceptions: Insights of the Typographic Vulnerability in Large Vision-Language Model

Abstract
Large Vision-Language Models (LVLMs) rely on vision encoders and Large Language Models (LLMs) to exhibit remarkable capabilities on various multi-modal tasks in the joint space of vision and language. However, the Typographic Attack, which disrupts Vision-Language Models (VLMs) such as Contrastive Language-Image Pretraining (CLIP), has also been expected to be a security threat to LVLMs. Firstly, we verify typographic attacks on current well-known commercial and open-source LVLMs and uncover the widespread existence of this threat. Secondly, to better assess this vulnerability, we propose the most comprehensive and largest-scale Typographic Dataset to date. The Typographic Dataset not only considers the evaluation of typographic attacks under various multi-modal tasks but also evaluates the effects of typographic attacks, influenced by texts generated with diverse factors.Based on the evaluation results, we investigate the causes why typographic attacks may impact VLMs and LVLMs, leading to three highly insightful discoveries. By the examination of our discoveries and experimental validation in Typographic Dataset, we reduce the performance degradation from to when LVLMs confront typographic attacks.
1 Introduction
The maturation of Large Language Models (LLMs) [3, 6, 8, 33, 37, 42, 42] in conjunction with the advancements in pre-trained vision encoders [35, 22] has catalyzed a growing interest in the domain of Large Vision-Language Models (LVLMs) of processing vision-language tasks [2, 23, 32, 7, 27, 49] and has expansive potential applications [13, 43]. However, the introduction of multi-modalities entails the emergence of supplementary security challenges, primarily attributable to the fact that each additional modality brings more vulnerabilities, which adversaries might potentially exploit [31, 12, 28].
Pre-trained vision encoders of VLMs, such as Contrastive Language-Image Pretraining (CLIP) [35] encounter a distinctive form of attack known as the typographic attack. Recent studies [4, 31, 5] show that typographic attacks can impair their zero-shot classification capabilities, deliberately introducing typographic errors into images to mislead CLIP. They claim that the underlying reason for typographic threats could be the shift in model attention before and after the introduction of typography. Therefore, LLaVA [27, 26], which is a typical type of LVLMs, is expected to inherit similar typographic weakness when incorporating the vision encoder CLIP. Moreover, as illustrated in Fig. 1, this weakness is prevalent across the majority of LVLMs. The possible reason for the typographic susceptibility in LVLMs might be inferred as a similar cause for this phenomenon occurring in CLIP. However, the mechanism for typographic effects in LVLMs remains an unexplored domain.
In this work, we systematically explore the distractibility of LVLMs by typography and try to uncover the actual reasons why LVLMs could be easily affected by typographic attacks. To measure the distractibility of LVLMs by typography, we introduce a specialized Typographic Dataset (TypoD), which serves as a Visual Question Answering (VQA) challenge, is derived from datasets including ImageNet [11], DAQUAR [30], CountBench [34], and A-OKVQA [40]. The Typographic Dataset meticulously assesses the distractibility of LVLMs by typography across perception-oriented and cognition-oriented tasks, including object recognition, visual attribute detection, enumeration, and commonsense reasoning as shown in Fig. 2. The Creation of TypoD could be separated into two stages, (1) Factor Exploring Stage: We explore typographic factors that may influence the effects of typographic attacks, encompassing font size, opacity, color attributes, and spatial positioning of typography within images. It is observed that a conspicuous positive correlation is evident between the visibility of typography and the efficacy of distraction. (2) Factor Fixing Stage: Through exploration of the typographic factors that can produce the most effective attacks above, in this stage, we fix the selected typography and generate two sub-type datasets of different scales, named TypoD-Base and TypoD-Large, each containing 1570 and 20000 images respectively. The TypoD-Base and TypoD-Large will provide a robust data platform for comprehensively evaluating the distractibility by typographic attacks and exploring the underlying cause of such threats.
To explore the vulnerability of LVLMs within the TypoD, We employ two typical LVLMs, including LLaVA and InstructBLIP [10]. Our empirical study reveals a substantive decrement in performance when applying LVLMs in the TypoD. In order to explore the authentic reason why typographic attacks occur, we compare the responses of the VLMs and LVLMs using the same vision encoder structure to the same typographic image. we observe that LVLMs could somewhat discriminate visual contents and typography in the images. However, the corresponding VLMs would fail to correctly differentiate between image content and added typography. Inspired by this phenomenon, we reveal that by augmenting the input information of the text-encoder in VLMs, particularly through offering more specific options to match the information from the image modality, VLMs can effectively differentiate between the image content and the inserted typography. Additionally, through Grad-CAM visualization [41], we observe that aside from the attention drawn by the addition of typographic text in the vision encoder, the amount of information contained in the options provided for the text-encoder also contributes to a certain degree of attention transfer.
Based on this observation, we uncover that by expanding input prompts’ information for LLMs within LVLMs, a similar phenomenon of attention transfer occurs towards visual contents of the image subjected to typographic attacks. Moreover, through inspecting attention maps of sequences in LVLMs, we further discover that increasing the demand for image descriptions in the prompt facilitates the generation of crucial information of visual contents of the image, which aids LVLMs in approaching the correct answers in VQA tasks. Therefore, in this scenario, the question part in the prompt not only extensively queries the image modality but also simultaneously queries the newly generated text modality information and interacts with newly generated crucial information, thus further preventing LVLMs from being excessively influenced by the added typographic texts when addressing the corresponding questions. Based on this, we present the following discoveries:
Discoveries: (1) The reason for occurring typographic attacks on VLMs and LVLMs lies in the attraction of attention caused by the inserted typographic text; (2) When VLMs and LVLMs confront typographic attacks, influences on the attention of vision encoders stem not only from direct modifications to the image but also from the guidance of text modality; (3) In multi-modal tasks, when confronting typographic attacks, VLMs and LVLMs necessitate cross-modal attention matching rather than solely on information from uni-modality to approach the final correct answers.
Our Contributions: In summary, the contributions of this work are manifold and may be encapsulated as follows:
We introduce the Typographic Dataset, which is the current largest scale platform to assess how typography can compromise the problem-solving capacities of LVLMs across various multi-modal tasks and typographic factors.
In our study, we have initially completed the most comprehensive and largest-scale evaluation of typographic attack performance under LVLMs to date.
Through exhaustive experiments and analysis, we present three intrinsic discoveries to elucidate the underlying reasons for typographic vulnerability in VLMs and LVLMs. This paves the authentic direction for future research on typographic attacks in multi-modal models.
Based on our discoveries, we narrow the average performance gap of LVLMs across different tasks under the scale of Typographic Dataset, whether the typography is added or not, to . Compared to the initial performance loss of experienced by LVLMs when subjected to the typographic attack, this paper has achieved a decrease of .
2 Related Work
2.1 Large Multimodal Models
The LLMs recently undergo a paradigmatic transformation. Instead of just working with text, these models can understand and respond to visual information, leading to the emergence of LVLMs. Integrating multiple modalities has marked this evolution, notably the visual modality [45] by incorporating pre-trained vision encoders and LLMs. Many architectures have been advanced to forge and refine the synthesis between text and other modalities. Notable among these are methodologies [23] that employ learnable queries to distill visual information, subsequently aligning the generation of language from LLMs upon visual features. Architectures such as MiniGPT-4 and LLaVA [49, 27, 26] have adopted a projection layer that effectively reconciles visual features derived from pre-trained vision encoders with the textual embeddings inherent to LLMs. In a distinct approach, Img2LLM [18] leverage frozen vision encoders and LLMs, deploying them modularly to facilitate zero-shot visual question answering without requiring additional fine-tuning. A suite of benchmarks [14, 44, 20] has been meticulously evaluated, affirming LVLMs demand acute visual perception and comprehensive understanding.
2.2 Vision Encoder
Vision encoders constitute an essential component of LVLMs, serving the critical feature extraction function from visual inputs. These encoders distill salient information from images and transmute this data into a high-dimensional embedding space. Among the vision encoders available, the CLIP model [35] stands as a paragon. CLIP is trained using a contrastive learning objective, which aims to align the representations of images and their corresponding text captions [38, 39, 19] in the same embedding space. This means that during training, these vision encoders learn to associate images with text captions so that the corresponding pairs are closer to the embedding space than non-corresponding pairs. This intrinsic capacity to bridge the semantic gap between visual and linguistic modalities underpins their versatility across a spectrum of downstream tasks, which span from the realm of zero-shot image classification [9, 15, 47] to the domain of semantic segmentation [48, 36].
2.3 Typographic Attack
The robustness of the CLIP against a specific form of attack, namely typographic attacks, has been questioned. Recent works [16, 4] provide empirical evidence illustrating a significant vulnerability in CLIP’s zero-shot classification robustness, which elucidates that the presence of typographic text within images can mislead CLIP to an erroneous classification.. The core issue arises from CLIP’s propensity to prioritize text within images when deriving a classification, a characteristic that can be exploited by embedding misleading typographic text, thereby inducing a misclassification [31, 5]. This susceptibility is notably troubling given that such attacks involve minor alterations to visual inputs, which is sufficient to derail the performance of CLIP.
3 Typographic Dataset
This section introduces the TypoD creation process, testing models, instruction design, and various typographic factors that could affect distractibility.
3.1 Dataset Setting
For a dataset particularly designed to demonstrate the distractibility of typographic attacks, different task categories, instruction designs, and typographic factors are the settings that need to be explored before creating.
Dataset Tasks: Perception and cognition are the core functionalities of LVLMs, and our evaluation centers on how typography influences LVLMs across four perception and cognition subtasks. Perception and cognition tasks could reflect the ability of authentic content interpretation and logical analysis for LVLMs. Concretely, TypoD encompass four multi-modal tasks, (1)Object Recognition: By leveraging the ImageNet [11], containing millions of labeled images spanning 1,000 object categories, we randomly choose 500 and 5000 images as normal images and their corresponding categories as ground truths from ImageNet. We also randomly select a different category for each image as a typo and construct a typographic image by printing the typo into the normal image. (2)Visual Attribute Detection: We use DAQUAR [30] and Visual7W [50] as our base datasets, which consists of images and associated questions and answers about the objects in the images and their attributes. Specifically, we choose 190 and 5000 questions about the colors of particular objects, their corresponding images as normal images, and their answers as ground truths. For every question, we introduce a randomly chosen incorrect color as a typo and create a typographic image by printing the typo into the normal image. (3)Enumeration: To test whether LVLMs would miscount the number of objects due to the influence of typographic text, We utilize the CountBench [34] and TallyQA[1], consisting of images that display a varying amount of items. We select the 390 and 5000 images as normal images, and their corresponding object counts as ground truths. For each image, we randomly select a different number as a typography and construct a typographic image. (4) Commonsense Reasoning: LVLMs answer questions in commonsense reasoning using various world knowledge. We leverage the A-OKVQA [40, 25], focusing on questions requiring world knowledge to answer. In particular, we select 500 and 5000 questions, their corresponding images as normal images, and answers as ground truths. For each question, we randomly choose a different choice from their related multiple choices as a typography and construct a typographic image by printing it into the normal image.
TypoD Scale | Factor Exploring | Factor Fixing | |||||
WTypo | FS | FO | FC | FP | TypoD-B | TypoD-L | |
Object | 5000 | 2500 | 2500 | 11500 | 8000 | 500 | 5000 |
Attribute | 5000 | 950 | 950 | 4370 | 3040 | 190 | 5000 |
Enumeration | 5000 | 1900 | 1900 | 8740 | 6080 | 380 | 5000 |
Reasoning | 5000 | 2500 | 2500 | 11500 | 8000 | 500 | 5000 |
Overall | 20000 | 7850 | 7850 | 36110 | 25120 | 1570 | 20000 |
Instruction Design: To quantify the distractibility of LVLMs by typography, our instruction is to let the model answer the correct option directly. Consequently, our instructional design is bifurcated into two distinct components: a task-related question and an explicit prompt enjoining the model to ”Answer with the option’s letter from the given choices directly”. This instruction paradigm is strategically crafted to delimit output to the specified options, thus facilitating the evaluation of accuracy metrics.
Typographic Factors: Various typographic factors could affect what a typo looks like in an image. Here, we mainly focus on font size, opacity, color attributes, and spatial positioning of typographic text, as demonstrated in Fig. 3. In particular, we set five size levels for font size spanning from 3 pixels to 15 pixels, increasing by three at each level. Regarding opacity, we put five opacity levels for font opacity spanning from 25% to 100%, the latter being completely opaque. To test LVLMs’ sensitivity to different font colors of typographic text, we select eight representative colors with their light and dark shades from spectrum colors, amounting to twenty-four colors. Regarding the spatial positioning of typographic text, we divide an image into a grid of four rows and four columns, resulting in sixteen equal-sized sections. We place typographic text in these sections to assess LVLMs’ sensitivity to typographic text in different positions.
3.2 Dataset Details
In this subsection, we depict the details for our TypoD, including dataset scale and particular generating process. We first select 1570 images for Object Recognition, Visual Attribute, Enumeration and Reasoning tasks as the Base (B) scale of our TypoD. Consequently, according to the describing and generating methods of typographic factors in Sec. 3.1, the first stage of creating TypoD is Factor Exploring. We need to alter different typographic factors to generate datasets of a certain scale to explore the influence of typographic patterns with different factors on LVLMs’ performance. Specifically, in each subtask, based on the initially selected 1570 image instances Without Typography (WTypo), We expand TypoD by incorporating additional variations, including 5 types each of Font Size (FS) and Font Opacity (FO), along with 23 types of Font Color (FC) and 16 types of Font Position (FP). Therefore, in the factor exploring stage, the scales of our TypoD could attain and instances respectively for WTypo and factors expansion (FS, FO, FC and FP). Sequentially, we would fix the typographic pattern with the factor that could generate the most powerful attacking effect and apply this kind of typography to the WTypo for generating our Typographic Dataset Base (TypoD-B). Furthermore, for the potentially more comprehensive typographic validation, we expand the TypoD-B to TypoD Large (TypoD-L) scale according to the selecting methods of image instances in Sec. 3.1. The WTypo-B is also expanded to WTypo-Large (WTypo-L) with the same scale as TypoD-L. WTypo-B and WTypo-L form the final WTypo dataset. The specific statistics of the TypoD scale is presented in Table 1.
4 Exploration of Typographic Attacks
LVLMs and VLMs utilize the similar vision encoder structure to acquire visual information. Consequently, some characteristic exploring conclusion targeting VLMs are straightforwardly extended to LVLMs. Previous works [4, 31, 5] indicate that typographic attacks pose significant risks to VLMs, attributing their impact to the substantial diversion of attention in the vision encoder of VLMs through simply adding typographic text. Hence, our initial motivation for studying typography under LVLMs, as well as accounting for the reasons behind typographic distraction in LVLMs, also stem from the related statements in those previous works. However, through exploration within this subsection, we uncover deeper reasons for typographic threats emerging under LVLMs, originating from attention grabbing by typography. Furthermore, alongside uncovering the corresponding truths, we have also accomplished the suppression of typographic attacks in both VLMs and LVLMs. Based on our uncovering, we pave the way for genuine exploration in the future research targeting typographic attacks under both vision-language tasks. In this subsection, we choose CLIP and LLaVA as the typical VLMs and LVLMs to execute our exploration.
Misconceptions about Vision Encoder: As stated above, the reason for typographic attacks in CLIP can be attribute to the fact that after adding typographic text to original images, the vision encoder used by CLIP would be guided by this text, directing its attention solely to the corresponding added regions. Therefore, since the same vision encoder structure as CLIP is employed for acquiring visual information, it can be intuitively assumed that LLaVA is susceptible to typographic attacks and the occurring reason would also be similar as CLIP. Furthermore, it can be inferred that an image causing typographic distraction in CLIP would have a similar effect when applied to LLaVA.
According to Fig.4, we observe that LLaVA could somewhat discriminate visual contents and typographic text in the images. When employing the exact same vision encoder as CLIP in LLaVA to match text information, CLIP failed to correctly differentiate between image content and typography. This suggests (1) LLaVA can give correct responses to typographical queries, indicating information extraction ability of concurrently used vision encoder is intact; (2) CLIP’s inability to provide accurate answers may stem from flawed guidance in its information extraction process, specifically, incomplete semantic information provided by the text options for the input of text encoder in CLIP. Inspired by such findings, we discover that the reason why CLIP is susceptible to typographic attacks in zero-shot classification is the limited semantic text options that can be chosen, like classifying an image by adopting simple zero-shot text options “an image of {dog, cat}”. Then, we prove that CLIP could almost avoid typographic attacks in zero-shot classification if more informative prompts are provided, as demonstrated in Fig. 4 (a). By providing additional information to the prompt text, such as including more options like ”an image of a {dog, cat} with the word ’dog’ written on it or an image of the word {dog, cat} ”, CLIP can correctly distinguish between the visual content and the added typographic elements under a typographic attack. This demonstrates that CLIP itself is not flawed and its vision encoder can accurately extracts complete information. The crucial issue lies in the varying levels of information contained in the text options provided for the text encoder.
To gain a deeper understanding of this phenomenon, we utilize Grad-CAM [41] to visualize the specific attention regions within the vision encoder of CLIP. In the Appendix, we observe the modifications in attention areas before and after adding typographic text to multiple images, thereby validating the claim from previous work that the primary attention of VLMs is attracted by typographic text. Subsequently, in Fig. 4 (a), we investigate the variations in the attention of CLIP after sequentially employing different text options from Fig. 4 (b), where brighter regions indicate stronger attention received in those areas. Through observing the 6 Grad-CAM images in Fig. 4 (b), we discover that (1) by comparing image pairs (2-3/5-6), the primary attention of the word ”cat” in the corresponding text options concentrate on the original content of the original image, namely the facial area of the cat, while the word ”dog” directs more attention to the added typographic ”Dog” text region. (2) By comparing image pairs (2-5/3-6), when the semantics in the text options shift from specifically referring to ”image” toward ”word”, the attention of the vision encoder would shift to the region of added typographic text, consequently reducing its attention on the original image content. (3) By comparing images (1,2,4,6) collectively, when the text option includes the word ”cat”, the attention of CLIP on the image content, specifically the facial region of the cat, is greater than that on the added typographic ”Dog” word. Conversely, the opposite is observed when the word ”dog” exists in the text option.
These above observations demonstrate that the vision encoder of CLIP has effectively understood the semantics of the words ”cat” and ”dog” in the text options, directing attention to the cat’s facial area, symbolizing the original image content, and to the added typographic ”dog” text, respectively. Furthermore, we propose a discovery that the semantic differences and the amount of information contained in the provided text input options significantly affect the attention of vision encoder in CLIP. And this variation in attention further becomes a key factor in determining the occurrence of typographic threats when using CLIP, and even overall VLMs. Thus we could obtain Discovery 1 mentioned above.
Multi-modal Query Capabilities: Currently, there is a consensus in vision-language tasks that when subjecting LVLMs to robustness attacks and interference, the security threat posed by simultaneously attacking multiple modalities is significantly greater than that of disrupting an uni-modality [29, 24, 17, 46]. The main reason for this phenomenon is the inherent asymmetry in information content between different modalities in multi-modal interactions. When a low informative modality queries information from a high informative modality, it leads to a deficiency in querying capability. In specific vision-language tasks, the text modality would exhibit a deficient capability when querying information from the visual modality. Building upon this weakness, Co-Attack [46] and SGA [29] achieve transferable black-box attacks across different VLMs. Similarly, FigStep [17] accomplishes multi-modal jailbreak attacks, a feat unattainable under the uni-modality conditions.
Through the description above , we observe that in CLIP, the addition of typographic text would result in direct attention distraction in the vision encoder. Additionally, the enhancement of input information for the text encoder could lead the visual attention of VLMs towards the overall image content. Therefore, based on this analysis and the claim in the previous works, the core reason for the emergence of typographic threat lies in the fact that the vision encoder directs excessive attention toward the added typographic text. The attentional attracting ability of typographic text reduces the overall focus of the vision encoder on the other general image content, resulting in incorrect responses. As LLaVA employs the same vision encoder structure as CLIP, the factors influencing the variation in visual attention should be the same as the factors in CLIP. The underlying cause of typographic vulnerability in LLaVA is likely akin to that observed in CLIP.
However, compared to the simple vision-language information matching when using CLIP, LLaVA, with the incorporation of Vicuna, achieves a richer multi-modal interaction in vision-language settings by more informative language prompts. If LLaVA experiences typographic attacks, it is for the same reason as CLIP, the added typographic text attracts the attention of the vision encoder. Based on the three discoveries above, we can suppress the typographic threat by augmenting the prompt information in LLaVA, thereby reducing the attention of the vision encoder on the typographic text region.
Specifically, this involves enhancing the semantics of image descriptions in the language prompt, aiming to strip as much image-specific information as possible and present it as direct textual information. This augmentation increases the amount of information in the language modality during multi-modal interactions. Moreover, this modality dissociation operation effectively addresses the issue of insufficient querying capability mentioned above in cross-modal scenarios. Fig. 5 (b) illustrates that LLaVA focuses more intently on the visual elements of the image (with the red areas indicating its focal points) when presented with an informative prompt. In contrast, it primarily focuses on typographic errors when responding to a simple prompt. Moreover, as indicated in Fig. 5 (c), the prompt, belonging to the language modality, not only queries the original image content but can also utilize newly generated language responses as query objects (light areas indicate higher levels of attention from LLaVA and corresponding tokens with high attention values to the final answer are highlighted in the chatbox). This significantly enhances the querying capability of the language modality. Thus Discovery 2 and Discovery 3 could be summarized above.
5 Experiments
In this section, we examine the distractibility of LVLMs on TypoD in both Factor Exploring Stage and Factor Fixing Stage. To understand how typography can distract LVLMs, we investigate the role of vision encoders and prompt LVLMs to articulate their thought process when confronting typographic text. To scrutinize LVLMs’ susceptibility to diversions in the Typographic Dataset, we mainly test on two popular LVLMs, including LLaVA [27, 26] and InstructBLIP [10, 21].
Tasks | TypoD-B(%) | TypoD-L(%) | ||||||||||
LLaVA-v1.5 | InstructBLIP | LLaVA-v1.5 | InstructBLIP | |||||||||
ACC | ACC- | GAP | ACC | ACC- | GAP | ACC | ACC- | GAP | ACC | ACC- | GAP | |
Obj | 97.8 | 35.6 | 62.2 | 97.8 | 66.4 | 31.4 | 97.9 | 45.4 | 52.5 | 97.9 | 65.6 | 32.3 |
Vis | 89.5 | 59.5 | 30.0 | 86.8 | 59.5 | 27.3 | 89.2 | 72.0 | 17.2 | 79.0 | 61.7 | 17.3 |
Enu | 74.4 | 40.0 | 34.4 | 84.2 | 58.4 | 25.8 | 88.6 | 62.1 | 26.5 | 85.6 | 39.3 | 46.3 |
Rea | 88.3 | 45.7 | 42.6 | 83.3 | 59.4 | 23.9 | 94.8 | 54.1 | 40.7 | 84.6 | 58.1 | 26.5 |
Overall | 87.3 | 45.2 | 42.3 | 88.0 | 60.9 | 27.1 | 82.3 | 49.9 | 32.4 | 75.7 | 47.2 | 28.5 |
Evaluation on TypoD: In the beginning, through adopting the data of the Factor Exploring Stage in TypoD, we explore various typographic factors that could affect distractibility, including Font Size, Font Opacity, Spatial Positioning and Font Color for added typographic text in Object Recognition (Obj), Visual Attribute (Vis), Enumeration (Enu) as well as Reasoning (Rea) sub-tasks. In this section, we present the the performance of object recognition, which best represents the VQA task, under various typical typographic factors, as illustrated in Fig. 6 (a). More detailed results with more various factors of other sub-tasks are show in Appendix Sec. 7.
Regarding Font Size and Opacity, Fig. 6 (a) shows a conspicuous positive correlation between the visibility of typographic text and the efficacy of typographic attacks. Furthermore, less visible typographic text still result in a significant reduction in models’ accuracy. typographic text with a font size of 6 pixels and an opacity of 20% can still cause an 11% and 16.6% drop in accuracy, respectively. Concerning Spatial Positioning and Font Color, Fig. 6 (a) shows that typographic text, regardless of their color or position, can significantly impact the accuracy of models. The phenomenon of typographic attacks to remain effective despite variations in color and location may be utilized to craft images laden with typographic text that, while seemingly ordinary, contain inconspicuous errors that can mislead LVLMs. This can be achieved by selecting a font color that merges with the background, making it hard to spot, and positioning it in an area that tends not to draw the eye.
According to the exploration results of typographic factors in Fig. 6 and results in Appendix Sec. 7, we fix the final typographic factors as {15px, 100%, R2C2 and White}. Using this setting, the sub-datasets, namely TypoD-B and TypoD-L, in the Factor Fixing stage of our TypoD are introduced. Next, we validate the performance degradation of LVLMs on Obj, Vis, Enu and Rea in Typo-D and Typo-L. In Typo-B, compared to the performance of LLaVA-v1.5 and InstructBLIP in WTypo-B, there would occur and performances drop. In Typo-L, the performances drop compared to WTypo-L are and respectively. We observe that typographic attacks indeed significantly degrade the performance of LVLMs. However, there are differences in the robustness of different LVLMs models against this threat, with InstructBLIP outperforming LLaVA-v1.5. For a more detailed exploration, we utilized LLaVA-v1.5 in our exploratory phase of LVLMs.
Tasks | Prompt 1 | Prompt 2.1 | Prompt 2.2 | Prompt 2.3 | Prompt 3 | ||||||||||
ACC | ACC- | GAP | ACC | ACC- | GAP | ACC | ACC- | GAP | ACC | ACC- | GAP | ACC | ACC- | GAP | |
Obj | 97.8 | 58.8 | 39.0 | 98.0 | 71.6 | 26.4 | 98.8 | 75.4 | 23.4 | 99.0 | 76.4 | 22.6 | 98.4 | 79.0 | 19.4 |
Vis | 83.5 | 68.7 | 14.8 | 93.8 | 83.0 | 10.7 | 93.3 | 85.1 | 8.2 | 92.8 | 84.5 | 8.28 | 93.8 | 84.5 | 9.2 |
Enu | 84.2 | 54.4 | 29.7 | 90.7 | 67.3 | 23.4 | 91.5 | 70.0 | 21.5 | 92.1 | 72.6 | 19.4 | 92.3 | 72.1 | 20.2 |
Rea | 89.1 | 46.0 | 43.0 | 91.3 | 55.3 | 35.9 | 90.4 | 58.5 | 31.8 | 87.3 | 58.8 | 28.4 | 84.6 | 60.3 | 24.2 |
Overall | 88.6 | 57.0 | 31.6 | 93.4 | 69.3 | 24.1 | 93.5 | 72.2 | 21.2 | 92.8 | 73.1 | 19.7 | 92.3 | 74.0 | 18.2 |
Exploring Discoveries: In order to further validate the discoveries in Fig. 4, we initially examine more image instances in the Appendix Sec. 8 to observe the similarity score in vision-language modal comparison and verify the distract effects of visual attention in Grad-CAM images. In addition, to assess the impact of changes in input information of the text encoder of CLIP on the severity of typographic threats in the zero-shot classification task, we randomly select 500 images from ImageNet and create their typographic versions by printing a random class as a typographic error in images. Hence we get a set of normal images and a set of typographic images. Then we craft image-matching texts for these images using Set 1 and Set 2 templates. Set 1: [an image of {label}, an image of {typo}]. Set 2: [an image of {label} with a word {typo} written on top of it, an image of {typo} with a word {typo} written on top of it, an image of {label}, an image of {typo}]. Through separate testing on CLIP-Base (CLIP-B) and CLIP-Large (CLIP-L), according to the Fig. 6, adopting the more informative Set 2 as templates would result in the enhancements of and compared to the less informative Set 1 in the zero-shot classification task.
For validating the dicoveries in Fig. 5, we need to augment prompts with additional semantics related to ”describing image content” to obtain new responses in text language modality. The text response generated by this new prompt, according to Discovery 2, on one hand, could mitigate the visual attention already attracted by the typographic text due to the increased information in the text modality. Also according to Discovery 3, on the other hand, the additonal information in the text modality can also match attention with the prompt, further directing overall attention favorably towards the correct response. we specifically implement the semantics of the above prompts by: 1. Directly adding requirements for describing images to the prompt; 2. Using multi-step prompts to generate more text language responses. Exemplary instructed prompts are illustrated following.
Prompt 1: Focus on the visual aspects of the image, including colors, shapes, composition, and any notable visual themes. Answer with the option’s letter from the given choices directly.
Prompt 2: (1) Provide a description of the image to answer the following question; (2) Provide a detailed visual description of the image to answer the following question;(3)Focus on the visual aspects of the image, including colors, shapes, composition, and any notable visual themes. Provide a detailed visual description of the image to answer the following question.
Prompt 3: Focus on the visual aspects of the image, including colors, shapes, composition, and any notable visual themes. Provide a detailed visual description of the image to answer the following question. Then based on your previous description, please delve deeper into the visual details of the image and include any subtle details or elements that were not covered in your initial description to answer the following question.
Compared to the GAP of for LLaVA-v1.5 in TypoD-B as shown in Table 2, the three prompts we test in Table 3 result in reductions in GAP to , , and , respectively. This indicates that adding semantics related to image description in prompts is indeed effective, but there may be a threshold, which may arise from the inherent gap in querying between multi-modal information. Additionally, by comparing the GAP of , , and for Prompt 2.1, 2.2, and 2.3, it further demonstrates that through the incremental increase of semantics describing image content, there is a slight stair-step suppression effect on typographic threats. More results of various prompts are presented in Appendix Sec. 8.
Ablation Study: Except for the above four multi-modal sub-tasks, in the Appendix Sec. 9 we also explore the performance of typographic vulnerability in the textual image. Additionally, we also tested a more intuitive approach to prompt semantic modification for typographic attacks. As presented in Fig. 6 (c), a further GAP decrease could be attained by inserting ”unaffected by typographic text” in the above new Prompts.
6 Conclusion
In this paper, we conduct verification of typographic attacks on leading LVLMs and confirm their prevalent existence. We presented an extensive Typographic Dataset to delve deeper into this susceptibility, evaluating typographic risks across different multi-modal tasks and typographic factors. Our research reveals three principal findings regarding the effects of typographic attacks on LVLMs. Inspired by these findings, through experimental validation with the Typographic Dataset, we reduce performance degradation from to when LVLMs confront typographic attacks.
Supplementary Material
7 Evaluation on TypoD
7.1 Details of Evaluated LVLMs
We provide more detailed results when LVLMs confront images with typographic text (typo) in various TypoD sub-type datasets. All experiments are conducted on NVIDIA A40 GPUs, and we test four state-of-the-art open-source models, including LLaVA-v1.5, InstructBLIP, LLaVA-v1.6 and MiniGPT4-v2. The evaluated versions are detailed below:
Model | Vision Encoder/LLMs |
LLaVA-v1.5 | CLIP-L-336px, Vicuna-13B |
InstructBLIP | CLIP-L, Vicuna-7B |
LLaVA-v1.6 | CLIP-L-336px, Vicuna-13B |
LLaVA-v1.6 | CLIP-L-336px, Hermes-Yi-34B |
MiniGPT4-v2 | EVA-CLIP, LLama-2-7B |
7.2 More Quantitative Results
Table 5 illustrates the impact of a simple typographic text within images on the performance of LVLMs across four sub-tasks in TypoD-B that test perception and cognition. Compared to the GAP of for LLaVA-v1.5 as shown in Table 2, LLaVA-v1.6-13B, LLaVA-v1.6-34B and MiniGPT4-v2 we test in Table 5 result in narrower GAPs of , , and , respectively. It indicates that powerful LVLMs are less susceptible to disruption by typographic attacks, but they are still vulnerable to them.
Table 8 to Table 11 show the distractibility of LVLMs to typographic texts with various typographic factors across four distinct sub-tasks in TypoD-B. In particular, we set five size levels for font size spanning from 3 pixels to 15 pixels. Regarding opacity, we put five opacity levels for font opacity spanning from 25% to 100%, the latter being completely opaque. To test LVLMs’ sensitivity to different font colors of typos, we select eight representative colors with their light and dark shades (e.g., ’lred’ and ’dred’) from spectrum colors. Regarding the spatial positioning of typos, we divide an image into a grid of four rows and four columns, resulting in sixteen equal-sized sections (e.g., ’R1C1’ denotes a typographic error located at Row 1, Column 1.).
8 Exploring Discoveries
8.1 CLIP with Informative Text Encoder Inputs
CLIP could almost avoid typographic attacks in zero-shot classification if more informative prompts are provided, as demonstrated in Fig. 4. In this subsection, we provide more examples to illustrate above discovery as presented in Fig. 7. By providing additional information to the prompt text, such as including more options like ”an image of a {stingray, tape player} with the word ’tape player’ written on it, CLIP shifts its attention area and correctly distinguishes between the visual contents and the added typographic texts under a typographic attack. This demonstrates that CLIP itself is not flawed and its vision encoder can accurately extract complete information. The crucial issue lies in the varying levels of information contained in the text options provided for the text encoder.
8.2 LLaVA with Informative Prompts
Compared to the simple vision-language information matching when using CLIP, LLaVA, with the incorporation of Vicuna as LLMs, achieves a richer multi-modal interaction in vision-language fields through more informative prompts. LLaVA can suppress the typographic threat by augmenting the prompt information, thereby reducing the attention of the vision encoder on the typographic text region. Fig. 5 and Fig. 8 illustrates: based on the chat-box in (a), (b) shows that LLaVA focuses more intently on the visual elements of the image (with the red areas indicating its focal attention areas) when presented with an informative prompt. In contrast, it primarily focuses on typographic errors when responding to a simple prompt. Moreover, (c) shows that the informative prompts, belonging to the language modality, not only query the original image content but can also utilize newly generated language responses as query objects (light areas indicate higher levels of attention from LLaVA and corresponding tokens with high attention values to the final answer are highlighted in the chatbox). This significantly enhances the querying capability of the language modality and further alleviate the incorrect attention targeting to the typo area.
8.3 Performance of Different Informative Prompts in LLaVA-v1.6
For further validating the discoveries in Fig. 5, we evaluate informative prompts in Table 3 on LLaVA-v1.6, as shown in Table 6. Compared to the GAP of for LLaVA-v1.6 in TypoD-B, the five prompts we test result in narrower GAPs of , , , and , respectively. This indicates that adding semantics related to image description in prompts is indeed effective.
On the one hand, the text response generated by this new prompt could mitigate the visual attention already attracted by the typographic text due to the increased information in the text modality. On the other hand, the additional information in the text modality can also further direct models’ attention favorably toward the correct response.
9 Ablation Study
9.1 Impacts of Typographic Texts on Images Containing Texts
In order to evaluate how typographic texts affect the performance of LVLMs on images containing texts. We craft a typographic arithmetic dataset. In the arithmetic computation, LVLMs need to comprehend arithmetic expressions embedded in images. Here, we focus on relatively easy arithmetic problems like addition, subtraction, multiplication, and division. We randomly selected two numbers and chose an arithmetic operation to compose an arithmetic problem. The correct answer is assigned as the ground truth, and a random number is selected as a typo. Then both the arithmetic problem and the typographic text are printed into the image. The distractibility of LVLMs with different typographic factors in the arithmetic computation task is shown in Table 12.
9.2 Prompts Instructed to Ignore Typographic Texts
Additionally, we test a more intuitive approach to prompt semantic modification for typographic attacks on LLaVA-v1.5. We add the additional words ”unaffected by typographic texts.” to each Prompts mentioned in Table 3 respectively. Then we get new Prompt 1-1, Prompt 2-1 and Prompt 3-1 which are instructed to ignore typographic texts. As shown in Table 7, compared to GAPs of unmodified prompts in Table 3 which are , and respectively, their instructed versions of ignoring typographic texts result in narrower GAPs of , , . This suggests that incorporating instructions to overlook typographic texts within prompts can genuinely enhance performance.
Tasks | LLaVA-v1.6-13B | LLaVA-v1.6-34B | MiniGPT4-v2 | ||||||
ACC | ACC- | GAP | ACC | ACC- | GAP | ACC | ACC- | GAP | |
Object Recognition | 98.0 | 44.0 | 54.0 | 99.18 | 77.15 | 22.03 | 94.8 | 86.6 | 8.21 |
Visual Attribute | 89.23 | 82.05 | 7.18 | 97.93 | 94.85 | 3.08 | 95.9 | 94.36 | 1.54 |
Enumeration | 70.53 | 44.21 | 26.32 | 95.25 | 79.21 | 16.04 | 75.26 | 70.0 | 5.26 |
Reasoning | 89.74 | 45.88 | 43.86 | 95.74 | 71.63 | 24.11 | 84.6 | 77.64 | 6.96 |
Overall | 86.87 | 54.03 | 32.84 | 97.025 | 80.71 | 16.31 | 87.63 | 82.14 | 5.49 |
Tasks | Prompt 1 | Prompt 2.1 | Prompt 2.2 | Prompt 2.3 | Prompt 3 | ||||||||||
ACC | ACC- | GAP | ACC | ACC- | GAP | ACC | ACC- | GAP | ACC | ACC- | GAP | ACC | ACC- | GAP | |
Object Recognition | 98.4 | 49.6 | 48.8 | 97.6 | 77.3 | 20.2 | 98.1 | 78.9 | 19.2 | 97.6 | 81.5 | 16.0 | 97.0 | 84.3 | 12.6 |
Visual Attribute | 92.3 | 82.0 | 10.2 | 97.9 | 94.8 | 3.1 | 97.9 | 95.8 | 2.0 | 96.9 | 95.3 | 1.5 | 97.4 | 95.3 | 2.0 |
Enumeration | 76.0 | 50.7 | 25.2 | 91.8 | 70.2 | 21.5 | 91.8 | 71.3 | 20.5 | 91.5 | 76.5 | 15.0 | 92.8 | 77.6 | 15.2 |
Reasoning | 90.3 | 44.8 | 45.4 | 89.5 | 60.0 | 29.4 | 86.2 | 63.6 | 22.5 | 86.1 | 67.6 | 18.4 | 83.2 | 67.6 | 15.5 |
Overall | 89.2 | 56.8 | 32.4 | 94.2 | 75.6 | 18.6 | 93.5 | 77.4 | 16.0 | 93.0 | 80.2 | 12.7 | 92.6 | 81.2 | 11.3 |
Tasks | Prompt 1-1 | Prompt 2-1 | Prompt 3-1 | ||||||
ACC | ACC- | GAP | ACC | ACC- | GAP | ACC | ACC- | GAP | |
Object Recognition | 97.80 | 69.00 | 28.80 | 98.60 | 85.40 | 13.20 | 98.20 | 86.87 | 11.32 |
Visual Attribute | 85.64 | 75.38 | 10.26 | 97.33 | 90.67 | 6.66 | 96.26 | 90.67 | 5.59 |
Enumeration | 75.79 | 34.47 | 41.32 | 92.37 | 72.63 | 19.74 | 91.58 | 72.89 | 18.69 |
Reasoning | 89.54 | 49.50 | 40.04 | 83.77 | 60.59 | 23.18 | 81.89 | 61.88 | 20.01 |
Overall | 87.10 | 57.08 | 30.10 | 93.01 | 77.32 | 15.69 | 91.98 | 78.07 | 13.90 |
Factors | Settings | LLaVA-v1.5 | InstructBLIP |
ACC | ACC | ||
Font Size | 3px | 81.20 | 97.60 |
6px | 61.00 | 97.40 | |
9px | 47.60 | 94.80 | |
12px | 42.60 | 72.80 | |
15px | 40.00 | 62.20 | |
Font Opacity | 20 % | 97.60 | 94.80 |
40 % | 86.80 | 88.07 | |
60 % | 50.80 | 80.80 | |
80 % | 40.20 | 69.80 | |
100 % | 39.60 | 61.60 | |
Font Color | red | 41.20 | 78.20 |
orange | 45.00 | 79.60 | |
yellow | 37.80 | 69.20 | |
green | 47.20 | 83.00 | |
cyan | 43.20 | 78.20 | |
blue | 43.00 | 76.60 | |
purple | 47.80 | 83.60 | |
dred | 44.40 | 75.00 | |
dorange | 58.80 | 84.20 | |
dyellow | 57.60 | 86.00 | |
dgreen | 53.00 | 82.60 | |
dcyan | 52.60 | 85.00 | |
dblue | 44.60 | 73.80 | |
dpurple | 46.00 | 78.00 | |
lred | 45.40 | 79.20 | |
lorange | 42.20 | 75.20 | |
lyellow | 37.80 | 65.60 | |
lgreen | 42.20 | 74.00 | |
lcyan | 41.00 | 72.40 | |
lblue | 50.60 | 82.40 | |
lpurple | 48.50 | 82.00 | |
white | 38.60 | 64.40 | |
black | 42.00 | 72.00 | |
Position | R1C1 | 62.20 | 68.20 |
R1C2 | 49.20 | 64.60 | |
R1C3 | 44.80 | 62.80 | |
R1C4 | 63.20 | 65.60 | |
R2C1 | 56.80 | 67.00 | |
R2C2 | 47.20 | 66.00 | |
R2C3 | 41.60 | 65.80 | |
R2C4 | 60.20 | 67.60 | |
R3C1 | 54.00 | 68.20 | |
R3C2 | 43.00 | 64.20 | |
R3C3 | 39.20 | 64.20 | |
R3C4 | 58.00 | 67.40 | |
R4C1 | 59.00 | 68.80 | |
R4C2 | 46.60 | 65.00 | |
R4C3 | 44.80 | 65.40 | |
R4C4 | 63.40 | 70.60 |
Factors | Settings | LLaVA-v1.5 | InstructBLIP |
ACC | ACC | ||
Font Size | 3px | 85.86 | 86.39 |
6px | 85.86 | 85.86 | |
9px | 67.54 | 87.96 | |
12px | 63.35 | 76.44 | |
15px | 63.87 | 58.64 | |
Font Opacity | 20 % | 69.47 | 76.32 |
40 % | 64.21 | 71.05 | |
60 % | 60.00 | 65.79 | |
80 % | 60.00 | 65.79 | |
100 % | 59.47 | 56.32 | |
Font Color | red | 61.14 | 66.84 |
orange | 59.07 | 69.95 | |
yellow | 58.55 | 60.10 | |
green | 62.18 | 71.50 | |
cyan | 59.07 | 69.43 | |
blue | 60.10 | 67.36 | |
purple | 58.55 | 64.25 | |
dred | 62.69 | 67.36 | |
dorange | 61.66 | 64.77 | |
dyellow | 58.55 | 68.91 | |
dgreen | 62.69 | 69.43 | |
dcyan | 61.14 | 69.95 | |
dblue | 60.10 | 67.36 | |
dpurple | 62.18 | 66.84 | |
lred | 60.10 | 66.84 | |
lorange | 59.59 | 62.69 | |
lyellow | 58.03 | 54.40 | |
lgreen | 59.07 | 64.25 | |
lcyan | 58.55 | 62.69 | |
lblue | 60.62 | 68.39 | |
lpurple | 60.62 | 66.84 | |
white | 58.55 | 55.44 | |
black | 65.28 | 65.80 | |
Position | R1C1 | 81.54 | 55.90 |
R1C2 | 75.38 | 57.95 | |
R1C3 | 74.87 | 56.92 | |
R1C4 | 84.62 | 49.74 | |
R2C1 | 81.54 | 51.79 | |
R2C2 | 71.79 | 53.85 | |
R2C3 | 71.28 | 56.41 | |
R2C4 | 82.56 | 51.79 | |
R3C1 | 80.00 | 50.77 | |
R3C2 | 63.08 | 54.87 | |
R3C3 | 61.03 | 50.77 | |
R3C4 | 82.05 | 50.77 | |
R4C1 | 81.54 | 49.23 | |
R4C2 | 60.00 | 49.74 | |
R4C3 | 56.92 | 45.64 | |
R4C4 | 84.10 | 48.72 |
Factors | Settings | LLaVA-v1.5 | InstructBLIP |
ACC | ACC | ||
Font Size | 3px | 73.68 | 83.42 |
6px | 72.37 | 82.11 | |
9px | 49.21 | 80.79 | |
12px | 38.42 | 64.47 | |
15px | 35.26 | 54.74 | |
Font Opacity | 20 % | 66.32 | 76.32 |
40 % | 56.84 | 71.92 | |
60 % | 47.63 | 66.84 | |
80 % | 39.74 | 58.68 | |
100 % | 35.26 | 52.11 | |
Font Color | red | 25.79 | 48.68 |
orange | 31.58 | 56.58 | |
yellow | 26.32 | 53.16 | |
green | 34.47 | 58.42 | |
cyan | 28.95 | 55.79 | |
blue | 25.26 | 45.00 | |
purple | 31.58 | 50.53 | |
dred | 26.58 | 44.47 | |
dorange | 38.68 | 53.95 | |
dyellow | 39.47 | 57.63 | |
dgreen | 35.26 | 51.58 | |
dcyan | 35.53 | 54.21 | |
dblue | 26.58 | 41.32 | |
dpurple | 28.95 | 45.26 | |
lred | 35.79 | 58.42 | |
lorange | 41.05 | 59.21 | |
lyellow | 33.95 | 56.32 | |
lgreen | 37.63 | 59.74 | |
lcyan | 35.26 | 58.42 | |
lblue | 38.42 | 58.68 | |
lpurple | 39.21 | 58.16 | |
white | 38.68 | 54.21 | |
black | 24.21 | 37.11 | |
Position | R1C1 | 45.79 | 53.68 |
R1C2 | 46.32 | 51.58 | |
R1C3 | 43.95 | 49.74 | |
R1C4 | 46.58 | 51.84 | |
R2C1 | 40.00 | 52.11 | |
R2C2 | 39.47 | 52.63 | |
R2C3 | 40.53 | 50.00 | |
R2C4 | 40.53 | 51.05 | |
R3C1 | 41.58 | 50.26 | |
R3C2 | 37.89 | 48.95 | |
R3C3 | 37.37 | 49.21 | |
R3C4 | 38.68 | 49.47 | |
R4C1 | 44.21 | 45.79 | |
R4C2 | 40.53 | 45.00 | |
R4C3 | 40.53 | 46.84 | |
R4C4 | 43.68 | 48.42 |
Factors | Settings | LLaVA-v1.5 | InstructBLIP |
ACC | ACC | ||
Font Size | 3px | 87.73 | 84.71 |
6px | 83.30 | 84.31 | |
9px | 56.74 | 81.89 | |
12px | 51.91 | 64.59 | |
15px | 48.49 | 54.12 | |
Font Opacity | 20 % | 78.87 | 80.48 |
40 % | 64.19 | 74.65 | |
60 % | 56.94 | 67.00 | |
80 % | 51.91 | 57.75 | |
100 % | 48.49 | 52.11 | |
Font Color | red | 50.40 | 67.60 |
orange | 54.80 | 69.60 | |
yellow | 48.40 | 58.40 | |
green | 57.80 | 72.80 | |
cyan | 50.40 | 64.60 | |
blue | 53.80 | 65.60 | |
purple | 56.20 | 71.20 | |
dred | 57.20 | 65.80 | |
dorange | 64.40 | 74.60 | |
dyellow | 61.20 | 75.40 | |
dgreen | 60.20 | 73.80 | |
dcyan | 59.00 | 74.20 | |
dblue | 55.40 | 63.20 | |
dpurple | 59.00 | 68.60 | |
lred | 55.00 | 68.80 | |
lorange | 53.40 | 62.20 | |
lyellow | 49.80 | 55.20 | |
lgreen | 53.40 | 63.80 | |
lcyan | 51.80 | 60.00 | |
lblue | 58.40 | 71.00 | |
lpurple | 55.00 | 69.40 | |
white | 49.00 | 54.20 | |
black | 54.60 | 63.60 | |
Position | R1C1 | 62.37 | 52.60 |
R1C2 | 55.73 | 52.80 | |
R1C3 | 52.72 | 50.80 | |
R1C4 | 61.97 | 52.20 | |
R2C1 | 56.74 | 49.00 | |
R2C2 | 49.90 | 54.40 | |
R2C3 | 48.29 | 53.60 | |
R2C4 | 56.94 | 51.00 | |
R3C1 | 55.33 | 52.00 | |
R3C2 | 47.69 | 52.80 | |
R3C3 | 47.48 | 53.20 | |
R3C4 | 56.54 | 50.60 | |
R4C1 | 56.34 | 51.40 | |
R4C2 | 48.29 | 48.40 | |
R4C3 | 48.49 | 47.40 | |
R4C4 | 62.37 | 50.40 |
Factors | Settings | LLaVA-v1.5 | InstructBLIP |
ACC | ACC | ||
Font Size | 3px | 43.89 | 30.26 |
6px | 29.80 | 31.86 | |
9px | 18.60 | 23.05 | |
12px | 15.60 | 14.40 | |
15px | 13.80 | 12.40 | |
Font Opacity | 20 % | 15.20 | 8.40 |
40 % | 13.60 | 8.00 | |
60 % | 12.20 | 8.60 | |
80 % | 12.40 | 10.60 | |
100 % | 12.40 | 13.00 | |
Font Color | red | 11.20 | 12.45 |
orange | 15.20 | 14.66 | |
yellow | 19.60 | 12.45 | |
green | 14.60 | 13.86 | |
cyan | 14.00 | 10.64 | |
blue | 12.80 | 14.03 | |
purple | 11.40 | 12.63 | |
dred | 14.40 | 13.25 | |
dorange | 15.40 | 14.03 | |
dyellow | 13.00 | 14.46 | |
dgreen | 11.80 | 14.23 | |
dcyan | 13.20 | 13.86 | |
dblue | 14.00 | 14.83 | |
dpurple | 10.20 | 15.43 | |
lred | 14.40 | 12.05 | |
lorange | 16.60 | 13.43 | |
lyellow | 23.60 | 17.64 | |
lgreen | 18.40 | 16.63 | |
lcyan | 15.80 | 11.82 | |
lblue | 14.80 | 11.42 | |
lpurple | 13.80 | 11.82 | |
white | 40.20 | 28.11 | |
black | 12.80 | 14.43 | |
Position | R1C1 | 41.00 | 10.80 |
R1C2 | 21.60 | 12.80 | |
R1C3 | 16.20 | 11.00 | |
R1C4 | 32.20 | 7.20 | |
R2C1 | 27.40 | 9.40 | |
R2C2 | 9.80 | 15.2 | |
R2C3 | 11.40 | 9.20 | |
R2C4 | 22.40 | 4.80 | |
R3C1 | 30.40 | 7.00 | |
R3C2 | 14.20 | 9.20 | |
R3C3 | 16.80 | 10.40 | |
R3C4 | 36.80 | 7.80 | |
R4C1 | 36.20 | 10.80 | |
R4C2 | 20.00 | 14.20 | |
R4C3 | 21.40 | 13.80 | |
R4C4 | 37.80 | 11.80 |

### v3

@@
-Unveiling Typographic Deceptions: Insights of the Typographic Vulnerability in Large Vision-Language Model
+Unveiling Typographic Deceptions: Insights of the Typographic Vulnerability in Large Vision-Language Models
 Abstract
 Large Vision-Language Models (LVLMs) rely on vision encoders and Large Language Models (LLMs) to exhibit remarkable capabilities on various multi-modal tasks in the joint space of vision and language.
-However, the Typographic Attack, which disrupts Vision-Language Models (VLMs) such as Contrastive Language-Image Pretraining (CLIP), has also been expected to be a security threat to LVLMs.
+However, typographic attacks, which disrupt Vision-Language Models (VLMs) such as Contrastive Language-Image Pretraining (CLIP), have also been expected to be a security threat to LVLMs.
 Firstly, we verify typographic attacks on current well-known commercial and open-source LVLMs and uncover the widespread existence of this threat.
 Secondly, to better assess this vulnerability, we propose the most comprehensive and largest-scale Typographic Dataset to date.
-The Typographic Dataset not only considers the evaluation of typographic attacks under various multi-modal tasks but also evaluates the effects of typographic attacks, influenced by texts generated with diverse factors.Based on the evaluation results, we investigate the causes why typographic attacks may impact VLMs and LVLMs, leading to three highly insightful discoveries.
-By the examination of our discoveries and experimental validation in Typographic Dataset, we reduce the performance degradation from to when LVLMs confront typographic attacks.
-1 Introduction
-The maturation of Large Language Models (LLMs) [3, 6, 8, 33, 37, 42, 42] in conjunction with the advancements in pre-trained vision encoders [35, 22] has catalyzed a growing interest in the domain of Large Vision-Language Models (LVLMs) of processing vision-language tasks [2, 23, 32, 7, 27, 49] and has expansive potential applications [13, 43].
-However, the introduction of multi-modalities entails the emergence of supplementary security challenges, primarily attributable to the fact that each additional modality brings more vulnerabilities, which adversaries might potentially exploit [31, 12, 28].
-Pre-trained vision encoders of VLMs, such as Contrastive Language-Image Pretraining (CLIP) [35] encounter a distinctive form of attack known as the typographic attack.
+The Typographic Dataset not only considers the evaluation of typographic attacks under various multi-modal tasks but also evaluates the effects of typographic attacks, influenced by texts generated with diverse factors.
+Based on the evaluation results, we investigate the causes why typographic attacks impacting VLMs and LVLMs, leading to three highly insightful discoveries.
+During the process of further validating the rationality of our discoveries, we can reduce the performance degradation caused by typographic attacks from 42.07% to 13.90%.
+Code and Dataset are available in https://github.com/ChaduCheng/TypoDeceptions.
+Keywords:
+Vision-Language Model Typographic Attack Attention1 Introduction
+The maturation of Large Language Models (LLMs) [13, 41, 46, 51, 51] in conjunction with the advancements in pre-trained vision encoders [43, 31] has catalyzed a growing interest in the domain of Large Vision-Language Models (LVLMs) of processing vision-language tasks [2, 30, 35, 62] and has expansive potential applications [12, 53].
+However, the introduction of multi-modalities entails the emergence of supplementary security challenges, primarily attributable to the fact that each additional modality brings more vulnerabilities, which adversaries might potentially exploit [38, 25, 5, 36, 52, 16, 27, 17, 21].
+Pre-trained vision encoders of VLMs, such as Contrastive Language-Image Pretraining (CLIP) [43], encounter a distinctive form of attack known as the typographic attack.
 Recent studies [3, 40, 4] show that typographic attacks can impair their zero-shot classification capabilities, deliberately introducing typographic errors into images to mislead CLIP.
-They claim that the underlying reason for typographic threats could be the shift in model attention before and after the introduction of typography.
-Therefore, LLaVA [27, 26], which is a typical type of LVLMs, is expected to inherit similar typographic weakness when incorporating the vision encoder CLIP.
-Moreover, as illustrated in Fig.
-1, this weakness is prevalent across the majority of LVLMs.
-The possible reason for the typographic susceptibility in LVLMs might be inferred as a similar cause for this phenomenon occurring in CLIP.
-However, the mechanism for typographic effects in LVLMs remains an unexplored domain.
+They claim that the underlying reason for typographic threats could be deficiencies in vision encoder’s ability to capture visual information, and the shift in model attention [23, 45] before and after the introduction of typography.
+Therefore, LLaVA [35, 34], which is a typical type of LVLMs, is assumed to inherit similar typographic weakness when incorporating the same vision encoder as CLIP.
+As illustrated in Fig.
+1, this weakness is indeed prevalent across the majority of LVLMs.
+And the possible reason for the typographic susceptibility in LVLMs might be inferred as a similar cause for this phenomenon occurring in CLIP.
+However, the true mechanism for typographic effects in LVLMs remains an unexplored domain.
 In this work, we systematically explore the distractibility of LVLMs by typography and try to uncover the actual reasons why LVLMs could be easily affected by typographic attacks.
-To measure the distractibility of LVLMs by typography, we introduce a specialized Typographic Dataset (TypoD), which serves as a Visual Question Answering (VQA) challenge, is derived from datasets including ImageNet [11], DAQUAR [30], CountBench [34], and A-OKVQA [40].
-The Typographic Dataset meticulously assesses the distractibility of LVLMs by typography across perception-oriented and cognition-oriented tasks, including object recognition, visual attribute detection, enumeration, and commonsense reasoning as shown in Fig.
+To measure the distractibility of LVLMs by typography, we introduce a specialized Typographic Dataset (TypoD), which serves as a Visual Question Answering (VQA) challenge under typographic attacks.
+TypoD is derived from datasets including ImageNet [11], DAQUAR [39], CountBench [42], and A-OKVQA [49].
+The TypoD meticulously assesses the distractibility of LVLMs by typography across perception-oriented and cognition-oriented tasks, including object recognition, visual attribute detection, enumeration, and commonsense reasoning as shown in Fig.
 2.
@@
 Additionally, through Grad-CAM visualization [50], we observe that aside from the attention drawn by the addition of typographic text in the vision encoder, the amount of information contained in the options provided for the text-encoder also contributes to a certain degree of attention transfer.
-Based on this observation, we uncover that by expanding input prompts’ information for LLMs within LVLMs, a similar phenomenon of attention transfer occurs towards visual contents of the image subjected to typographic attacks.
+Based on this observation, we uncover that by expanding information of input prompts for LLMs within LVLMs, a similar phenomenon of attention transfer occurs towards visual contents of the image subjected to typographic attacks.
 Moreover, through inspecting attention maps of sequences in LVLMs, we further discover that increasing the demand for image descriptions in the prompt facilitates the generation of crucial information of visual contents of the image, which aids LVLMs in approaching the correct answers in VQA tasks.
@@
 Discoveries: (1) The reason for occurring typographic attacks on VLMs and LVLMs lies in the attraction of attention caused by the inserted typographic text; (2) When VLMs and LVLMs confront typographic attacks, influences on the attention of vision encoders stem not only from direct modifications to the image but also from the guidance of text modality; (3) In multi-modal tasks, when confronting typographic attacks, VLMs and LVLMs necessitate cross-modal attention matching rather than solely on information from uni-modality to approach the final correct answers.
-Our Contributions: In summary, the contributions of this work are manifold and may be encapsulated as follows:
-We introduce the Typographic Dataset, which is the current largest scale platform to assess how typography can compromise the problem-solving capacities of LVLMs across various multi-modal tasks and typographic factors.
+Our Contributions are manifold and may be encapsulated as follows:
+We introduce the Typographic Dataset (TypoD), which is the current largest platform to assess how typography can compromise the problem-solving capacities of LVLMs across various multi-modal tasks and typographic factors.
 In our study, we have initially completed the most comprehensive and largest-scale evaluation of typographic attack performance under LVLMs to date.
@@
 2 Related Work
-2.1 Large Multimodal Models
-The LLMs recently undergo a paradigmatic transformation.
-Instead of just working with text, these models can understand and respond to visual information, leading to the emergence of LVLMs.
+Large Vision-Language Model (LVLMs): The LLMs recently underwent a paradigmatic transformation.
+Instead of just working with text, LVLMs can understand and respond to visual information.
 Integrating multiple modalities has marked this evolution, notably the visual modality [58] by incorporating pre-trained vision encoders and LLMs.
+Thus, LVLMs achieve excellent performance in numerous traditional vision tasks [57, 6, 56, 55, 7, 14, 8].
 Many architectures have been advanced to forge and refine the synthesis between text and other modalities.
@@
 A suite of benchmarks [15, 54, 28] has been meticulously evaluated, affirming LVLMs demand acute visual perception and comprehensive understanding.
-2.2 Vision Encoder
-Vision encoders constitute an essential component of LVLMs, serving the critical feature extraction function from visual inputs.
+Vision Encoder: Vision encoders constitute an essential component of LVLMs, serving the critical feature extraction function from visual inputs.
 These encoders distill salient information from images and transmute this data into a high-dimensional embedding space.
@@
 This intrinsic capacity to bridge the semantic gap between visual and linguistic modalities underpins their versatility across a spectrum of downstream tasks, which span from the realm of zero-shot image classification [9, 18, 60] to the domain of semantic segmentation [61, 44].
-2.3 Typographic Attack
-The robustness of the CLIP against a specific form of attack, namely typographic attacks, has been questioned.
-Recent works [16, 4] provide empirical evidence illustrating a significant vulnerability in CLIP’s zero-shot classification robustness, which elucidates that the presence of typographic text within images can mislead CLIP to an erroneous classification..
-The core issue arises from CLIP’s propensity to prioritize text within images when deriving a classification, a characteristic that can be exploited by embedding misleading typographic text, thereby inducing a misclassification [31, 5].
+Typographic Attack: The robustness of the CLIP against a specific form of attack, namely typographic attacks, has been questioned.
+Recent works [19, 3] provide empirical evidence illustrating a significant vulnerability in CLIP’s zero-shot classification robustness, which elucidates that the presence of typographic text within images can mislead CLIP to an erroneous classification.
+The core issue arises from propensity of CLIP to prioritize text within images when deriving a classification, a characteristic that can be exploited by embedding misleading typographic text, thereby inducing a misclassification [40, 4].
 This susceptibility is notably troubling given that such attacks involve minor alterations to visual inputs, which is sufficient to derail the performance of CLIP.
 3 Typographic Dataset
-This section introduces the TypoD creation process, testing models, instruction design, and various typographic factors that could affect distractibility.
+This section introduces the Typographic Dataset (TypoD) creation process.
 3.1 Dataset Setting
-For a dataset particularly designed to demonstrate the distractibility of typographic attacks, different task categories, instruction designs, and typographic factors are the settings that need to be explored before creating.
-Dataset Tasks: Perception and cognition are the core functionalities of LVLMs, and our evaluation centers on how typography influences LVLMs across four perception and cognition subtasks.
+For a dataset particularly designed to demonstrate the distractibility of typographic attacks, different task categories, instruction designs, and typographic factors are the settings that need to be explored before creating the dataset.
+Dataset Tasks: Perception and cognition are the core functionalities of LVLMs, and our evaluation centers on how typography influences LVLMs across four subtasks.
 Perception and cognition tasks could reflect the ability of authentic content interpretation and logical analysis for LVLMs.
-Concretely, TypoD encompass four multi-modal tasks, (1)Object Recognition: By leveraging the ImageNet [11], containing millions of labeled images spanning 1,000 object categories, we randomly choose 500 and 5000 images as normal images and their corresponding categories as ground truths from ImageNet.
+Concretely, TypoD encompasses four multi-modal tasks, (1)Object Recognition: By leveraging the ImageNet [11], containing millions of labeled images spanning 1,000 object categories, we randomly choose 500 and 5000 images as normal images and their corresponding categories as ground truths from ImageNet.
 We also randomly select a different category for each image as a typo and construct a typographic image by printing the typo into the normal image.
@@
 For each question, we randomly choose a different choice from their related multiple choices as a typography and construct a typographic image by printing it into the normal image.
-TypoD Scale | Factor Exploring | Factor Fixing | |||||
-WTypo | FS | FO | FC | FP | TypoD-B | TypoD-L | |
-Object | 5000 | 2500 | 2500 | 11500 | 8000 | 500 | 5000 |
-Attribute | 5000 | 950 | 950 | 4370 | 3040 | 190 | 5000 |
-Enumeration | 5000 | 1900 | 1900 | 8740 | 6080 | 380 | 5000 |
-Reasoning | 5000 | 2500 | 2500 | 11500 | 8000 | 500 | 5000 |
-Overall | 20000 | 7850 | 7850 | 36110 | 25120 | 1570 | 20000 |
 Instruction Design: To quantify the distractibility of LVLMs by typography, our instruction is to let the model answer the correct option directly.
-Consequently, our instructional design is bifurcated into two distinct components: a task-related question and an explicit prompt enjoining the model to ”Answer with the option’s letter from the given choices directly”.
+Consequently, our instructional design is bifurcated into two distinct components: a task-related question and an explicit prompt enjoining the model to "Answer with the option’s letter from the given choices directly".
 This instruction paradigm is strategically crafted to delimit output to the specified options, thus facilitating the evaluation of accuracy metrics.
@@
 We need to alter different typographic factors to generate datasets of a certain scale to explore the influence of typographic patterns with different factors on LVLMs’ performance.
-Specifically, in each subtask, based on the initially selected 1570 image instances Without Typography (WTypo), We expand TypoD by incorporating additional variations, including 5 types each of Font Size (FS) and Font Opacity (FO), along with 23 types of Font Color (FC) and 16 types of Font Position (FP).
-Therefore, in the factor exploring stage, the scales of our TypoD could attain and instances respectively for WTypo and factors expansion (FS, FO, FC and FP).
-Sequentially, we would fix the typographic pattern with the factor that could generate the most powerful attacking effect and apply this kind of typography to the WTypo for generating our Typographic Dataset Base (TypoD-B).
-Furthermore, for the potentially more comprehensive typographic validation, we expand the TypoD-B to TypoD Large (TypoD-L) scale according to the selecting methods of image instances in Sec.
+Specifically, in each subtask, based on the initially selected 1570 images Without Typography (WTypo), we expand TypoD by incorporating additional variations, including 5 types each of Font Size (FS) and Font Opacity (FO), along with 23 types of Font Color (FC) and 16 types of Position (P).
+Therefore, in the factor exploring stage, the scales of our TypoD could attain and instances respectively for WTypo and factors expansion (FS, FO, FC and P).
+Sequentially, we would fix the typographic pattern with the factor that could generate the most powerful attacking effect and apply this kind of typography to the WTypo Base (WTypo-B) for generating our Typographic Dataset Base (TypoD-B).
+Furthermore, for the potentially more comprehensive typographic validation, we expand the TypoD-B and WTypoD-B to TypoD Large (TypoD-L) and WTypo-Large (WTypo-L) scale according to the selecting methods of image instances in Sec.
 3.1.
-The WTypo-B is also expanded to WTypo-Large (WTypo-L) with the same scale as TypoD-L.
-WTypo-B and WTypo-L form the final WTypo dataset.
-The specific statistics of the TypoD scale is presented in Table 1.
+WTypoD-L could also be directly abbreviated as WTypoD.
+The specific statistics of the WTypoD, FS, FO, FC, FP, TypoD-B and TypoD-L scale are presented in Table 1.
+TypoD Scale | Factor Exploring | Factor Fixing | |||||
+WTypo | FS | FO | FC | FP | TypoD-B | TypoD-L | |
+Object | 5000 | 2500 | 2500 | 11500 | 8000 | 500 | 5000 |
+Attribute | 5000 | 950 | 950 | 4370 | 3040 | 190 | 5000 |
+Enumeration | 5000 | 1900 | 1900 | 8740 | 6080 | 380 | 5000 |
+Reasoning | 5000 | 2500 | 2500 | 11500 | 8000 | 500 | 5000 |
+Overall | 20000 | 7850 | 7850 | 36110 | 25120 | 1570 | 20000 |
 4 Exploration of Typographic Attacks
-LVLMs and VLMs utilize the similar vision encoder structure to acquire visual information.
-Consequently, some characteristic exploring conclusion targeting VLMs are straightforwardly extended to LVLMs.
-Previous works [4, 31, 5] indicate that typographic attacks pose significant risks to VLMs, attributing their impact to the substantial diversion of attention in the vision encoder of VLMs through simply adding typographic text.
-Hence, our initial motivation for studying typography under LVLMs, as well as accounting for the reasons behind typographic distraction in LVLMs, also stem from the related statements in those previous works.
-However, through exploration within this subsection, we uncover deeper reasons for typographic threats emerging under LVLMs, originating from attention grabbing by typography.
+LVLMs and VLMs utilize a similar vision encoder structure to acquire visual information.
+Consequently, some characteristic exploring conclusions targeting VLMs are straightforwardly extended to LVLMs.
+Previous works [3, 40, 4] indicate that typographic attacks pose significant risks to VLMs, attributing their impact to the substantial diversion of attention [23, 45] in the vision encoder of VLMs through simply adding typographic text.
+Hence, our initial motivation for studying typography under LVLMs, as well as accounting for the reasons behind typographic distraction in LVLMs, also stems from the related statements in those previous works.
+However, through exploration within this subsection, we uncover deeper reasons for typographic threats emerging under LVLMs, originating from attention grabbed by typography.
 Furthermore, alongside uncovering the corresponding truths, we have also accomplished the suppression of typographic attacks in both VLMs and LVLMs.
@@
 In this subsection, we choose CLIP and LLaVA as the typical VLMs and LVLMs to execute our exploration.
-Misconceptions about Vision Encoder: As stated above, the reason for typographic attacks in CLIP can be attribute to the fact that after adding typographic text to original images, the vision encoder used by CLIP would be guided by this text, directing its attention solely to the corresponding added regions.
-Therefore, since the same vision encoder structure as CLIP is employed for acquiring visual information, it can be intuitively assumed that LLaVA is susceptible to typographic attacks and the occurring reason would also be similar as CLIP.
+Misconceptions about Vision Encoder: As stated above, the reason for typographic attacks in CLIP can be attributed to the fact that after adding typographic text to original images, the vision encoder used by CLIP would be guided by this text, directing its attention solely to the corresponding added regions.
+Therefore, since the same vision encoder structure as CLIP is employed for acquiring visual information, it can be intuitively assumed that LLaVA is susceptible to typographic attacks and the occurring reason would also be similar to CLIP.
 Furthermore, it can be inferred that an image causing typographic distraction in CLIP would have a similar effect when applied to LLaVA.
@@
 When employing the exact same vision encoder as CLIP in LLaVA to match text information, CLIP failed to correctly differentiate between image content and typography.
-This suggests (1) LLaVA can give correct responses to typographical queries, indicating information extraction ability of concurrently used vision encoder is intact; (2) CLIP’s inability to provide accurate answers may stem from flawed guidance in its information extraction process, specifically, incomplete semantic information provided by the text options for the input of text encoder in CLIP.
+This suggests (1) LLaVA can give correct responses to typographical queries, indicating its information extraction ability is intact.
+(2) CLIP’s inability to provide accurate answers may stem from flawed guidance in its information extraction process, specifically, incomplete semantic information provided by the text options for the input of text encoder in CLIP.
 Inspired by such findings, we discover that the reason why CLIP is susceptible to typographic attacks in zero-shot classification is the limited semantic text options that can be chosen, like classifying an image by adopting simple zero-shot text options “an image of {dog, cat}”.
@@
 4 (a).
-By providing additional information to the prompt text, such as including more options like ”an image of a {dog, cat} with the word ’dog’ written on it or an image of the word {dog, cat} ”, CLIP can correctly distinguish between the visual content and the added typographic elements under a typographic attack.
-This demonstrates that CLIP itself is not flawed and its vision encoder can accurately extracts complete information.
+By providing additional information to the prompt text [22], such as including more options like "an image of a {dog, cat} with the word ’dog’ written on it or an image of the word {dog, cat} ", CLIP can correctly distinguish between the visual content and the added typographic elements under a typographic attack.
+This demonstrates that CLIP itself is not flawed and its vision encoder can accurately extract complete information.
 The crucial issue lies in the varying levels of information contained in the text options provided for the text encoder.
-To gain a deeper understanding of this phenomenon, we utilize Grad-CAM [41] to visualize the specific attention regions within the vision encoder of CLIP.
-In the Appendix, we observe the modifications in attention areas before and after adding typographic text to multiple images, thereby validating the claim from previous work that the primary attention of VLMs is attracted by typographic text.
+To better understand this phenomenon, we utilize Grad-CAM [50] to visualize the specific attention regions within the vision encoder of CLIP.
 Subsequently, in Fig.
@@
 Through observing the 6 Grad-CAM images in Fig.
-4 (b), we discover that (1) by comparing image pairs (2-3/5-6), the primary attention of the word ”cat” in the corresponding text options concentrate on the original content of the original image, namely the facial area of the cat, while the word ”dog” directs more attention to the added typographic ”Dog” text region.
-(2) By comparing image pairs (2-5/3-6), when the semantics in the text options shift from specifically referring to ”image” toward ”word”, the attention of the vision encoder would shift to the region of added typographic text, consequently reducing its attention on the original image content.
-(3) By comparing images (1,2,4,6) collectively, when the text option includes the word ”cat”, the attention of CLIP on the image content, specifically the facial region of the cat, is greater than that on the added typographic ”Dog” word.
-Conversely, the opposite is observed when the word ”dog” exists in the text option.
-These above observations demonstrate that the vision encoder of CLIP has effectively understood the semantics of the words ”cat” and ”dog” in the text options, directing attention to the cat’s facial area, symbolizing the original image content, and to the added typographic ”dog” text, respectively.
-Furthermore, we propose a discovery that the semantic differences and the amount of information contained in the provided text input options significantly affect the attention of vision encoder in CLIP.
+4 (b), we discover that (1) by comparing image pairs (2-3/5-6), the primary attention of the word "cat" in the corresponding text options concentrate on the original content of the original image, namely the facial area of the cat, while the word "dog" directs more attention to the added typographic "Dog" text region.
+(2) By comparing image pairs (2-5/3-6), when the semantics in the text options shift from specifically referring to "image" toward "word", the attention of the vision encoder would shift to the region of added typographic text, consequently reducing its attention on the original image content.
+(3) By comparing images (1,2,4,6) collectively, when the text option includes the word "cat", the attention of CLIP on the image content, specifically the facial region of the cat, is greater than that on the added typographic "Dog" word.
+Conversely, the opposite is observed when the word "dog" exists in the text option.
+The above discussions demonstrate that the vision encoder of CLIP has effectively understood the semantics of the words "cat" and "dog" in the text options, directing attention to the cat’s facial area, symbolizing the original image content, and to the added typographic "dog" text, respectively.
+Furthermore, we propose a discovery that the semantic differences and the amount of information contained in the provided text input options significantly affect the attention of the vision encoder in CLIP.
 And this variation in attention further becomes a key factor in determining the occurrence of typographic threats when using CLIP, and even overall VLMs.
-Thus we could obtain Discovery 1 mentioned above.
 Multi-modal Query Capabilities: Currently, there is a consensus in vision-language tasks that when subjecting LVLMs to robustness attacks and interference, the security threat posed by simultaneously attacking multiple modalities is significantly greater than that of disrupting an uni-modality [37, 32, 20, 59].
@@
 The underlying cause of typographic vulnerability in LLaVA is likely akin to that observed in CLIP.
-However, compared to the simple vision-language information matching when using CLIP, LLaVA, with the incorporation of Vicuna, achieves a richer multi-modal interaction in vision-language settings by more informative language prompts.
+However, compared to the simple vision-language information matching when using CLIP, LLaVA, with the incorporation of Vicuna, achieves a richer multi-modal interaction in vision-language fields through more informative prompts [22].
 If LLaVA experiences typographic attacks, it is for the same reason as CLIP, the added typographic text attracts the attention of the vision encoder.
@@
 Fig.
-5 (b) illustrates that LLaVA focuses more intently on the visual elements of the image (with the red areas indicating its focal points) when presented with an informative prompt.
+5 (b) illustrates that LLaVA focuses more intently on the visual elements of the image (with the red areas indicating its focal areas) when presented with an informative prompt.
 In contrast, it primarily focuses on typographic errors when responding to a simple prompt.
@@
 This significantly enhances the querying capability of the language modality.
-Thus Discovery 2 and Discovery 3 could be summarized above.
 5 Experiments
@@
 Overall | 87.3 | 45.2 | 42.3 | 88.0 | 60.9 | 27.1 | 82.3 | 49.9 | 32.4 | 75.7 | 47.2 | 28.5 |
-Evaluation on TypoD: In the beginning, through adopting the data of the Factor Exploring Stage in TypoD, we explore various typographic factors that could affect distractibility, including Font Size, Font Opacity, Spatial Positioning and Font Color for added typographic text in Object Recognition (Obj), Visual Attribute (Vis), Enumeration (Enu) as well as Reasoning (Rea) sub-tasks.
-In this section, we present the the performance of object recognition, which best represents the VQA task, under various typical typographic factors, as illustrated in Fig.
+Evaluation on TypoD: By adopting the data of the Factor Exploring Stage in TypoD, we explore various typographic factors that could affect distractibility, including Font Size, Font Opacity, Spatial Positioning, and Font Color for added typographic text in Object Recognition (Obj), Visual Attribute (Vis), Enumeration (Enu) as well as Reasoning (Rea) sub-tasks.
+In this section, we present the performance of object recognition, which best represents the VQA task, under various typical typographic factors, as illustrated in Fig.
 6 (a).
-More detailed results with more various factors of other sub-tasks are show in Appendix Sec.
-7.
+More detailed results with more various factors of other sub-tasks are shown in Appendix Sec.
+0.A.
 Regarding Font Size and Opacity, Fig.
 6 (a) shows a conspicuous positive correlation between the visibility of typographic text and the efficacy of typographic attacks.
-Furthermore, less visible typographic text still result in a significant reduction in models’ accuracy.
+Furthermore, less visible typographic text still results in a significant reduction in models’ accuracy.
 typographic text with a font size of 6 pixels and an opacity of 20% can still cause an 11% and 16.6% drop in accuracy, respectively.
 Concerning Spatial Positioning and Font Color, Fig.
-6 (a) shows that typographic text, regardless of their color or position, can significantly impact the accuracy of models.
+6 (a) shows that typographic text, regardless of its color or position, can significantly impact the accuracy of models.
 The phenomenon of typographic attacks to remain effective despite variations in color and location may be utilized to craft images laden with typographic text that, while seemingly ordinary, contain inconspicuous errors that can mislead LVLMs.
@@
 6 and results in Appendix Sec.
-7, we fix the final typographic factors as {15px, 100%, R2C2 and White}.
+0.A, we fix the final typographic factors as {15px, 100%, R2C2 and White}.
 Using this setting, the sub-datasets, namely TypoD-B and TypoD-L, in the Factor Fixing stage of our TypoD are introduced.
-Next, we validate the performance degradation of LVLMs on Obj, Vis, Enu and Rea in Typo-D and Typo-L.
+Next, we validate the performance degradation of LVLMs on Obj, Vis, Enu, and Rea in Typo-D and Typo-L.
 In Typo-B, compared to the performance of LLaVA-v1.5 and InstructBLIP in WTypo-B, there would occur and performances drop.
@@
 We observe that typographic attacks indeed significantly degrade the performance of LVLMs.
-However, there are differences in the robustness of different LVLMs models against this threat, with InstructBLIP outperforming LLaVA-v1.5.
+However, there are differences in the robustness of different LVLMs against this threat, with InstructBLIP outperforming LLaVA-v1.5.
 For a more detailed exploration, we utilized LLaVA-v1.5 in our exploratory phase of LVLMs.
@@
 Exploring Discoveries: In order to further validate the discoveries in Fig.
-4, we initially examine more image instances in the Appendix Sec.
-8 to observe the similarity score in vision-language modal comparison and verify the distract effects of visual attention in Grad-CAM images.
+4, we initially examine more image instances in Appendix Sec.
+0.B to observe the similarity score in vision-language modal comparison and verify the distract effects of visual attention in Grad-CAM images.
 In addition, to assess the impact of changes in input information of the text encoder of CLIP on the severity of typographic threats in the zero-shot classification task, we randomly select 500 images from ImageNet and create their typographic versions by printing a random class as a typographic error in images.
@@
 6, adopting the more informative Set 2 as templates would result in the enhancements of and compared to the less informative Set 1 in the zero-shot classification task.
-For validating the dicoveries in Fig.
-5, we need to augment prompts with additional semantics related to ”describing image content” to obtain new responses in text language modality.
-The text response generated by this new prompt, according to Discovery 2, on one hand, could mitigate the visual attention already attracted by the typographic text due to the increased information in the text modality.
-Also according to Discovery 3, on the other hand, the additonal information in the text modality can also match attention with the prompt, further directing overall attention favorably towards the correct response.
-we specifically implement the semantics of the above prompts by: 1.
+For validating the discoveries in Fig.
+5, we need to augment prompts with additional semantics related to "describing image content" to obtain new responses in text language modality.
+The text response generated by this new prompt, according to Discovery 2, on the one hand, could mitigate the visual attention already attracted by the typographic text due to the increased information in the text modality.
+Also according to Discovery 3, on the other hand, the additional information in the text modality can also match attention with the prompt, further directing overall attention favorably towards the correct response.
+we specifically implement the semantics of the above prompts by 1.
 Directly adding requirements for describing images to the prompt; 2.
@@
 Additionally, by comparing the GAP of , , and for Prompt 2.1, 2.2, and 2.3, it further demonstrates that through the incremental increase of semantics describing image content, there is a slight stair-step suppression effect on typographic threats.
-More results of various prompts are presented in Appendix Sec.
-8.
-Ablation Study: Except for the above four multi-modal sub-tasks, in the Appendix Sec.
-9 we also explore the performance of typographic vulnerability in the textual image.
+More results, such as evaluation on various prompts and models, are presented in Appendix Sec.
+0.B.
+Ablation Study: Except for the above four multi-modal sub-tasks, in Appendix Sec.
+0.C we also explore the impact of typography on images containing texts.
 Additionally, we also tested a more intuitive approach to prompt semantic modification for typographic attacks.
 As presented in Fig.
-6 (c), a further GAP decrease could be attained by inserting ”unaffected by typographic text” in the above new Prompts.
+6 (c), a further GAP decrease could be attained by prompts instructed to ignore typography modified from the above informative prompts.
 6 Conclusion
@@
 We presented an extensive Typographic Dataset to delve deeper into this susceptibility, evaluating typographic risks across different multi-modal tasks and typographic factors.
-Our research reveals three principal findings regarding the effects of typographic attacks on LVLMs.
-Inspired by these findings, through experimental validation with the Typographic Dataset, we reduce performance degradation from to when LVLMs confront typographic attacks.
-Supplementary Material
-7 Evaluation on TypoD
-7.1 Details of Evaluated LVLMs
+Appendix 0.A Evaluation on TypoD
+0.A.1 Details of Evaluated LVLMs
 We provide more detailed results when LVLMs confront images with typographic text (typo) in various TypoD sub-type datasets.
@@
 MiniGPT4-v2 | EVA-CLIP, LLama-2-7B |
-7.2 More Quantitative Results
+0.A.2 More Quantitative Results
 Table 5 illustrates the impact of a simple typographic text within images on the performance of LVLMs across four sub-tasks in TypoD-B that test perception and cognition.
@@
 Regarding the spatial positioning of typos, we divide an image into a grid of four rows and four columns, resulting in sixteen equal-sized sections (e.g., ’R1C1’ denotes a typographic error located at Row 1, Column 1.).
-8 Exploring Discoveries
-8.1 CLIP with Informative Text Encoder Inputs
+Tasks | LLaVA-v1.6-13B | LLaVA-v1.6-34B | MiniGPT4-v2 | ||||||
+ACC | ACC- | GAP | ACC | ACC- | GAP | ACC | ACC- | GAP | |
+Object Recognition | 98.0 | 44.0 | 54.0 | 99.18 | 77.15 | 22.03 | 94.8 | 86.6 | 8.21 |
+Visual Attribute | 89.23 | 82.05 | 7.18 | 97.93 | 94.85 | 3.08 | 95.9 | 94.36 | 1.54 |
+Enumeration | 70.53 | 44.21 | 26.32 | 95.25 | 79.21 | 16.04 | 75.26 | 70.0 | 5.26 |
+Reasoning | 89.74 | 45.88 | 43.86 | 95.74 | 71.63 | 24.11 | 84.6 | 77.64 | 6.96 |
+Overall | 86.87 | 54.03 | 32.84 | 97.025 | 80.71 | 16.31 | 87.63 | 82.14 | 5.49 |
+Appendix 0.B Exploring Discoveries
+0.B.1 CLIP with Informative Text Encoder Inputs
 CLIP could almost avoid typographic attacks in zero-shot classification if more informative prompts are provided, as demonstrated in Fig.
@@
 7.
-By providing additional information to the prompt text, such as including more options like ”an image of a {stingray, tape player} with the word ’tape player’ written on it, CLIP shifts its attention area and correctly distinguishes between the visual contents and the added typographic texts under a typographic attack.
+By providing additional information to the prompt text, such as including more options like "an image of a {stingray, tape player} with the word ’tape player’ written on it, CLIP shifts its attention area and correctly distinguishes between the visual contents and the added typographic texts under a typographic attack.
 This demonstrates that CLIP itself is not flawed and its vision encoder can accurately extract complete information.
 The crucial issue lies in the varying levels of information contained in the text options provided for the text encoder.
-8.2 LLaVA with Informative Prompts
+0.B.2 LLaVA with Informative Prompts
 Compared to the simple vision-language information matching when using CLIP, LLaVA, with the incorporation of Vicuna as LLMs, achieves a richer multi-modal interaction in vision-language fields through more informative prompts.
@@
 This significantly enhances the querying capability of the language modality and further alleviate the incorrect attention targeting to the typo area.
-8.3 Performance of Different Informative Prompts in LLaVA-v1.6
+0.B.3 Performance of Different Informative Prompts in LLaVA-v1.6 and InstructBLIP
 For further validating the discoveries in Fig.
-5, we evaluate informative prompts in Table 3 on LLaVA-v1.6, as shown in Table 6.
+5, we evaluate informative prompts in Table 3 on LLaVA-v1.6 and InstructBLIP, as shown in Table 6 and Table 7.
 Compared to the GAP of for LLaVA-v1.6 in TypoD-B, the five prompts we test result in narrower GAPs of , , , and , respectively.
@@
 On the other hand, the additional information in the text modality can also further direct models’ attention favorably toward the correct response.
-9 Ablation Study
-9.1 Impacts of Typographic Texts on Images Containing Texts
-In order to evaluate how typographic texts affect the performance of LVLMs on images containing texts.
-We craft a typographic arithmetic dataset.
-In the arithmetic computation, LVLMs need to comprehend arithmetic expressions embedded in images.
-Here, we focus on relatively easy arithmetic problems like addition, subtraction, multiplication, and division.
-We randomly selected two numbers and chose an arithmetic operation to compose an arithmetic problem.
-The correct answer is assigned as the ground truth, and a random number is selected as a typo.
-Then both the arithmetic problem and the typographic text are printed into the image.
-The distractibility of LVLMs with different typographic factors in the arithmetic computation task is shown in Table 12.
-9.2 Prompts Instructed to Ignore Typographic Texts
-Additionally, we test a more intuitive approach to prompt semantic modification for typographic attacks on LLaVA-v1.5.
-We add the additional words ”unaffected by typographic texts.” to each Prompts mentioned in Table 3 respectively.
-Then we get new Prompt 1-1, Prompt 2-1 and Prompt 3-1 which are instructed to ignore typographic texts.
-As shown in Table 7, compared to GAPs of unmodified prompts in Table 3 which are , and respectively, their instructed versions of ignoring typographic texts result in narrower GAPs of , , .
-This suggests that incorporating instructions to overlook typographic texts within prompts can genuinely enhance performance.
-Tasks | LLaVA-v1.6-13B | LLaVA-v1.6-34B | MiniGPT4-v2 | ||||||
-ACC | ACC- | GAP | ACC | ACC- | GAP | ACC | ACC- | GAP | |
-Object Recognition | 98.0 | 44.0 | 54.0 | 99.18 | 77.15 | 22.03 | 94.8 | 86.6 | 8.21 |
-Visual Attribute | 89.23 | 82.05 | 7.18 | 97.93 | 94.85 | 3.08 | 95.9 | 94.36 | 1.54 |
-Enumeration | 70.53 | 44.21 | 26.32 | 95.25 | 79.21 | 16.04 | 75.26 | 70.0 | 5.26 |
-Reasoning | 89.74 | 45.88 | 43.86 | 95.74 | 71.63 | 24.11 | 84.6 | 77.64 | 6.96 |
-Overall | 86.87 | 54.03 | 32.84 | 97.025 | 80.71 | 16.31 | 87.63 | 82.14 | 5.49 |
 Tasks | Prompt 1 | Prompt 2.1 | Prompt 2.2 | Prompt 2.3 | Prompt 3 | ||||||||||
 ACC | ACC- | GAP | ACC | ACC- | GAP | ACC | ACC- | GAP | ACC | ACC- | GAP | ACC | ACC- | GAP | |
-Object Recognition | 98.4 | 49.6 | 48.8 | 97.6 | 77.3 | 20.2 | 98.1 | 78.9 | 19.2 | 97.6 | 81.5 | 16.0 | 97.0 | 84.3 | 12.6 |
-Visual Attribute | 92.3 | 82.0 | 10.2 | 97.9 | 94.8 | 3.1 | 97.9 | 95.8 | 2.0 | 96.9 | 95.3 | 1.5 | 97.4 | 95.3 | 2.0 |
-Enumeration | 76.0 | 50.7 | 25.2 | 91.8 | 70.2 | 21.5 | 91.8 | 71.3 | 20.5 | 91.5 | 76.5 | 15.0 | 92.8 | 77.6 | 15.2 |
-Reasoning | 90.3 | 44.8 | 45.4 | 89.5 | 60.0 | 29.4 | 86.2 | 63.6 | 22.5 | 86.1 | 67.6 | 18.4 | 83.2 | 67.6 | 15.5 |
+Obj | 98.4 | 49.6 | 48.8 | 97.6 | 77.3 | 20.2 | 98.1 | 78.9 | 19.2 | 97.6 | 81.5 | 16.0 | 97.0 | 84.3 | 12.6 |
+Vis | 92.3 | 82.0 | 10.2 | 97.9 | 94.8 | 3.1 | 97.9 | 95.8 | 2.0 | 96.9 | 95.3 | 1.5 | 97.4 | 95.3 | 2.0 |
+Enu | 76.0 | 50.7 | 25.2 | 91.8 | 70.2 | 21.5 | 91.8 | 71.3 | 20.5 | 91.5 | 76.5 | 15.0 | 92.8 | 77.6 | 15.2 |
+Rea | 90.3 | 44.8 | 45.4 | 89.5 | 60.0 | 29.4 | 86.2 | 63.6 | 22.5 | 86.1 | 67.6 | 18.4 | 83.2 | 67.6 | 15.5 |
 Overall | 89.2 | 56.8 | 32.4 | 94.2 | 75.6 | 18.6 | 93.5 | 77.4 | 16.0 | 93.0 | 80.2 | 12.7 | 92.6 | 81.2 | 11.3 |
+Tasks | Prompt 1 | Prompt 2 | Prompt 3 | ||||||
+ACC | ACC- | GAP | ACC | ACC- | GAP | ACC | ACC- | GAP | |
+Obj | 97.8 | 66.4 | 31.4 | 98.0 | 87.2 | 10.79 | 98.4 | 89.2 | 9.2 |
+Vis | 86.15 | 62.05 | 24.1 | 76.92 | 60.31 | 16.61 | 80.0 | 65.13 | 14.87 |
+Enu | 84.21 | 58.42 | 25.79 | 80.7 | 67.17 | 13.53 | 92.54 | 80.39 | 12.15 |
+Rea | 83.3 | 59.56 | 23.74 | 82.86 | 63.91 | 18.95 | 79.64 | 63.71 | 15.93 |
+Overall | 87.86 | 61.60 | 26.25 | 84.61 | 69.64 | 14.97 | 87.64 | 74.60 | 13.03 |
 Tasks | Prompt 1-1 | Prompt 2-1 | Prompt 3-1 | ||||||
@@
 Overall | 87.10 | 57.08 | 30.10 | 93.01 | 77.32 | 15.69 | 91.98 | 78.07 | 13.90 |
+Appendix 0.C Ablation Study
+0.C.1 Prompts Instructed to Ignore Typographic Texts
+Additionally, we test a more intuitive approach to prompt semantic modification for typographic attacks on LLaVA-v1.5.
+We add the additional words "unaffected by typographic texts." to each Prompts mentioned in Table 3 respectively.
+Then we get new Prompt 1-1, Prompt 2-1 and Prompt 3-1 which are instructed to ignore typographic texts.
+As shown in Table 8, compared to GAPs of unmodified prompts in Table 3 which are , and respectively, their instructed versions of ignoring typographic texts result in narrower GAPs of , , .
+This suggests that incorporating instructions to overlook typographic texts within prompts can genuinely enhance performance.
+0.C.2 Detailed Results of Typogrpahic Factors in differernt sub-tasks
+In order to further evaluate how typography affect the performance of LVLMs.
+We evaluate LLaVA-v1.5 and InstructBLIP in TypoD-B with more various typographic factors in four sub-tasks.
+The results of object recognition, visual attribute detection, enumeration, and commonsense reasoning are presented in Table 9, Table 10, Table 11 and Table 12.
+Additionally, we also craft a typographic arithmetic dataset.
+In the arithmetic computation, LVLMs need to comprehend arithmetic expressions embedded in images.
+Here, we focus on relatively easy arithmetic problems like addition, subtraction, multiplication, and division.
+We randomly selected two numbers and chose an arithmetic operation to compose an arithmetic problem.
+The correct answer is assigned as the ground truth, and a random number is selected as a typo.
+Then both the arithmetic problem and the typographic text are printed into the image.
+The distractibility of LVLMs with different typographic factors in the arithmetic computation task is shown in Table 13.
 Factors | Settings | LLaVA-v1.5 | InstructBLIP |